Calculate the Rf using the following equation:

Rf = distance pigment migrated (mm) / distance solvent migrated (mm)

Table 1: Part 1: Chromatography Data (15 points)

Solvent

Distance from Original Line to Solvent Front

Number of Bands

Rf Factor

Acetic Acid

9cm

1

?

Acetone

7.2cm

3

?

Mineral Oil

5cm

1

?

Water

8.8

0

?
Answer #1

You have to taken silvent which is moved to the farthest distance in the chromatography. In this case solvent is acetic acid.

Now you have to divide each solvent cover distance to solvent distance on chromatography.

Solvent Rf value
Acetic acid 9/9=1
Acetone 7.2/9= 0.8
Mineral oil 5/9= 0.55
Water 8.8/9= 0.91